# Who Pays for Fairness? Rethinking Recourse under Social Burden

### v1

@@ 3.2 The Burden of False Negatives
In Equation 2, we measure the expected recourse cost while accounting for the classifier’s predictions, providing a more precise notion of fairness. However, it focuses exclusively on the model’s predictions $f(\mathbf{x})$, without accounting for the true class label. As a result, it fails to distinguish between individuals who were incorrectly denied a positive outcome (i.e., truly positive but misclassified) and those who were correctly assigned a negative outcome. In this section, we take a step further by focusing on minimizing the expected cost associated with positive instances that the classifier incorrectly predicts as negative. This next step naturally leads to the concept of social burden. [^2] Some works on algorithmic recourse (e.g., Kavouras et al. [29]) employ the term burden to refer to the average cost itself. However, in an earlier work by Milli et al. [35], burden is defined as the average cost incurred by instances that have a true positive class label. In this paper, we adopt the latter definition.
Social burden measures the expected cost incurred by individuals who are already qualified (true positives) but receive an unfavorable prediction by a classifier $f$ [35]. The social burden is zero if the classifier positively classifies all qualified individuals without requiring changes to their non-sensitive attributes. Conversely, the social burden is greater than zero when qualified individuals are assigned a negative classification and forced to alter their features to be recognized as valid. Formally, we can define social burden as:
@@
Let $f$ be the classifier, $g_{f}$ the recourse algorithm, and $\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))$ the cost function. We denote the expected social burden for the sensitive group $\boldsymbol{s}$ as $\mathbb{E}_{P(\boldsymbol{X}=\boldsymbol{x}|S=\boldsymbol{s},Y=1)}[\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))]$. Then, the expected cost in terms of the classifier’s true positive rate can be defined as:
$\mathcal{B}_{f,g}^{\boldsymbol{s}}=\mathbb{E}[\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))]\cdot(1-P(f(\boldsymbol{x})=1\mid S=\boldsymbol{s},Y=1))$ (3)
with the expectation over $p(X\mid S=\boldsymbol{s},Y=1,f(\boldsymbol{x})=0)$.
Similarly to Equation 2, the instances obtaining a positive prediction do not count towards the overall burden.
Equation 3 underscores the importance of considering the predictive performance of the base classifier when evaluating algorithmic recourse, a factor often overlooked in the recourse literature but central to algorithmic fairness; especially, since numerous studies have shown that machine learning classifiers tend to exhibit unequal error rates (e.g., TPR) across sensitive groups [41]. When error rates differ across groups, the burden of recourse is also likely to be unevenly distributed, potentially amplifying existing disparities. However, it is important to note that satisfying equality of opportunity (i.e., parity in TPR) does not guarantee a small gap in social burden.
@@ 3.3 Gap-Based Metrics can Mask Unfairness
Unfairness in recourse can be quantified by measuring gaps in cost or social burden across sensitive groups (e.g, $\Delta\mathcal{C}_{f,g}=\max_{s,s^{\prime}\in\mathcal{S}}|\mathcal{C}^{s}_{f,g}-\mathcal{C}^{s^{\prime}}_{f,g}|$ or $\Delta\mathcal{B}_{f,g}=\max_{s,s^{\prime}\in\mathcal{S}}|\mathcal{B}^{s}_{f,g}-\mathcal{B}^{s^{\prime}}_{f,g}|$). According to Equations 2 and 3, these disparities stem from three main sources. For cost gaps, the causes include: (a) disparities in AR, as captured by statistical parity [14]; (b) differences in the cost function, which may arise from unequal opportunities or socio-economic conditions; and (c) variations in the distribution $P(X\mid S=\boldsymbol{s},f(\boldsymbol{x})=0)$, reflecting potential population skews. For social burden gaps, they arise from: (a) disparities in true positive rates, captured by equality of opportunity [23], (b) unequal cost functions, and (c) skews in the distribution of qualified individuals denied recourse (i.e., $P(X\mid S=\boldsymbol{s},Y=1,f(\boldsymbol{x})=0)$). Importantly, this illustrates that satisfying statistical parity (equality of opportunity) does not guarantee an equal distribution of recourse costs (social burden) across groups.
Although comparing the gap of statistical measures across groups is a common approach in the fairness literature, it can obscure meaningful underlying unfairness. In fairness for prediction, a well-known limitation of parity-based definitions is that parity can be achieved by degrading performance for the privileged group without improving outcomes for the unprivileged one [33, 13]. As a result, groups may exhibit similar expected values for a given metric, even when those values are far from optimal. In recourse, this issue is especially critical when evaluating social burden, which ideally should be as low as possible (preferably zero). A small gap in burden does not necessarily indicate fairness, as parity might be reached by increasing the burden for the privileged group rather than reducing it for the unprivileged one. However, in many real-world contexts, such as access to essential goods and services, sacrificing overall well-being is not acceptable. Therefore, fairness efforts should aim to reduce the burden for all groups, rather than simply equalizing it.
We propose drawing inspiration from approaches in the fairness-in-prediction literature that address this limitation and complement difference-based disparity measures with a Rawlsian (i.e., minimax) perspective [5]. For instance, in addition to evaluating the gap in social burden, we can examine the worst-group social burden defined as $\max_{\boldsymbol{s}\in\mathcal{S}}\mathcal{B}_{f,g}^{\boldsymbol{s}}$. This perspective’s benefits are twofold: (a) it helps identify the most vulnerable group, and (b) it serves as a base for the development of fairness-enhancing recourse frameworks that go beyond minimizing burden disparities, aiming instead to reduce the social burden for all groups.
@@
5:   $y^{i}\leftarrow f^{(t-1)}(\boldsymbol{x}^{i})$ // Predict label
6:   $b_{f,g}^{i}\leftarrow\delta(\boldsymbol{x}^{i},g_{f}(\boldsymbol{x}^{i}))\mathbb{I}\{y^{i}=1\}$ // Compute burden
7:   $\mathcal{Q}\leftarrow\mathcal{Q}\cup b_{f,g}^{i}$
8:  end for
9:  $f^{(t)}\leftarrow\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}\phi(i,\mathcal{Q})\cdot\ell(f(\boldsymbol{x}^{i}),y^{i})$
10:end for
11:return  $f^{(T)}$
4 MISOB: MInimax SOcial Burden
In this section, we introduce a simple yet effective iterative training procedure, MISOB (MInimax SOcial Burden), that empirically reduces social burden (Equation 3) without degrading predictive performance and without requiring access to sensitive group attributes during training or inference. The method is outlined in Algorithm 1. Given a pre-trained classifier (line 1) [^3] Empirically, we observed that pretraining a base classifier to achieve reasonable accuracy prior to applying MISOB contributes to a more stable optimization process., the key idea is to incorporate social burden estimates directly into the training loop. At each iteration, the current classifier is used to generate predictions (line 5), from which the social burden for each instance is computed (line 6), and saved within $\mathcal{Q}$ (line 7). The model is then updated by minimizing a (weighted) loss function $\ell(f(\boldsymbol{x}^{i}),y^{i})$ (line 9), where weights are determined by the estimated burden $\phi(i,\mathcal{Q})$. Specifically, we propose assigning weights to instances $\boldsymbol{x}^{i}$ based on the following scheme:
$\phi(i,\mathcal{Q})=\begin{cases}1+CN\dfrac{b_{f,g_{f}}^{i}}{\sum_{j\in|\mathcal{Q}|}b_{f,g_{f}}^{j}}&\text{if }B>0\\ 1&\text{if }B=0\end{cases}$ (4)
where $N$ is the dataset (or batch) size, $B=\sum_{j\in|\mathcal{Q}|}b_{f,g_{f}}^{j}$ is the total burden with $b_{f,g}^{i}=\delta(\boldsymbol{x}^{i},g_{f}(\boldsymbol{x}^{i}))\cdot\allowbreak\mathbb{I}\{y^{i}=1\}$ being the burden of instance $i$, and $C$ is a hyperparameter controlling the influence of the burden term. Under this weighting scheme, each instance $\boldsymbol{x}^{i}$ is assigned a weight proportional to its contribution to the total burden across the dataset (or batch). In the degenerate case, where the total burden is zero, all instances are assigned equal weights. This reweighting scheme amplifies the influence of high-burden instances during training, encouraging the classifier to prioritize decisions that reduce unnecessary or costly adjustments imposed on individuals under the current decision boundary.
The proposed framework offers several advantages that enhance its practicality and broad applicability. [^4] The runtime of Algorithm 1 is determined by three factors: (i) the number of retraining steps $T$, (ii) the dataset size $N$, and (iii) the run-time complexity of the recourse method $O(K)$. In practice, datasets are typically much larger than the number of retraining iterations, and if a simple recourse method is used, we generally have $K<<T<<N$. As a result, the overall computational complexity is $O(N^{3})$. Efficiency can be further improved through batching, parallelization, or by approximating recourse computations. First, the method is agnostic to both the base classifier and the chosen recourse strategy, allowing seamless integration with a wide range of existing approaches. It also operates without access to sensitive attributes during training, avoiding legal, ethical, and practical challenges associated with collecting or using such data. Further, MISOB not only improves deployability but also naturally addresses intersectional fairness. Because group definitions are not required a-priori, intersectional subgroups can be introduced post-hoc at evaluation time, enabling fairness assessments across multiple dimensions without retraining. This flexibility makes the framework particularly well-suited for real-world applications where fairness concerns span diverse and evolving group identities.
5 Empirical Evaluation
In this section, we evaluate the effectiveness of MISOB in reducing the social burden across sensitive groups without compromising the overall predictive performance of the classifier. [^5] We will release the source code on GitHub under a permissive license. To this end, we experiment with various datasets, base classifiers and recourse methods, assessing how integrating burden information into the training process improves fairness guarantees in the algorithmic pipeline under study.
5.1 Experimental Setting
Datasets.
We consider the real-world dataset Adult [3] with sensitive groups defined by different characterizations of the sensitive attributes (race and gender), considering also intersectional groups. For results with other datasets, refer to Appendix C.
Base model.
For the base model, we consider a neural network (NN) with 2 hidden layers composed of 128 neurons each. We have chosen this base classifier because it can be merged with multiple different recourse methods proposed in the literature. Experiments with additional base classifiers can be found in Appendix C. We combine the NN with MISOB as well as with the method to enhance prediction fairness in terms of equality of opportunity (i.e., parity in TPR) by Hardt et al. [23] (POSTPRO). We employ POSTPRO because it is compatible with any base classifier, enabling fair comparisons with MISOB under identical predictive models.
Recourse methods.
We consider three widely adopted recourse methods, selected for their benchmarking popularity, available implementations, and ease of integration in an agnostic setting: Growing Spheres (GS) [31], the method by Wachter et al. [49] (WT), and the Counterfactual Conditional Heterogeneous Variational Autoencoder (CCHVAE) by Pawelczyk et al. [37]. We consider the implementations and hyperparameter settings from the benchmark by Pawelczyk et al. [38].
We assess the effectiveness of the proposed framework using overall accuracy and a range of fairness metrics, covering the whole pipeline (prediction and recourse). Specifically, we evaluate the maximum disparity across groups and the worst-group values for social burden, TPR, cost and AR. Recourse cost is quantified using the $\ell_{2}$ distance ($\times 100$) between the original instance and the recourse target. Importantly, our method operates entirely without access to sensitive attributes. Note that, to compute fairness metrics, whether for prediction or recourse, the sensitive information of instances is required, an inherent limitation of fairness metrics themselves, as discussed in Buyl and De Bie [6]. For a more detailed explanation on the experimental setting refer to Appendix B.
5.2 Results
Table 2 presents the average results over 10 random train/test splits. The findings show that training the base classifier under the MISOB framework enhances fairness both throughout the whole recourse-aware algorithmic pipeline, as well as at the prediction level, across all groups, without sacrificing overall accuracy. Notably, MISOB enhances worst-case group outcomes and, in many cases, reduces disparities across groups, achieving fairness improvements without degrading the performance for privileged groups. These results also highlight an important limitation of gap-based fairness metrics: small gaps can mask universally poor outcomes across groups, giving a misleading sense of fairness. In contrast, MISOB can produce configurations with slightly larger gaps but better metric values for every group, thereby promoting greater overall well-being. For instance, in the first block of results in Table 2, we observe that under POSTPRO, both sensitive groups achieve an equal average TPR of 0.37. In contrast, with MISOB, the average TPRs increase to 0.52 and 0.63 for the two groups, respectively. Although the gap between groups is larger, both sensitive groups achieve higher and more desirable TPR values overall.
[…]
Recourse Fairness acc Burden (Eq( 3)) TPR Cost (Eq( 2)) AR
@@ 5.2 Results
Method strategy overall $\uparrow$} worst $\downarrow$} $\Delta$ $|\downarrow|$} worst $\uparrow$} $\Delta$ $|\downarrow|$} worst $\downarrow$} $\Delta$ $|\downarrow|$} worst $\uparrow$} $\Delta$ $|\downarrow|$}
race
GS - 0.81±0.02 4.56±0.01 0.03±0.02 0.27±0.03 0.08±0.08 115.69±1.19 28.37±1.95 0.06±0.01 0.06±0.02

### v3

@@ 3.2 The Burden of False Negatives
In Eqn. 2, we measure the expected recourse cost while accounting for the classifier’s predictions, providing a more precise notion of fairness. However, it focuses exclusively on the model’s predictions $f(x)$, without accounting for the true class label. As a result, it fails to distinguish between individuals who were incorrectly denied a positive outcome (i.e., truly positive but misclassified) and those who were correctly assigned a negative outcome. In this section, we take a step further by focusing on minimizing the expected cost associated with positive instances that the classifier incorrectly predicts as negative. This next step naturally leads to the concept of social burden. [^4] Some works on algorithmic recourse (e.g., kavouras2023fairness) employ the term burden to refer to the average cost itself. However, in an earlier work by milli2019social, burden is defined as the average cost incurred by instances that have a true positive class label. In this paper, we adopt the latter definition. Social burden measures the expected cost incurred by individuals who are already qualified (true positives) but receive an unfavorable prediction by a classifier $f$ (milli2019social). The social burden is zero if the classifier positively classifies all qualified individuals without requiring changes to their non-sensitive attributes. Conversely, the social burden is greater than zero when qualified individuals are assigned a negative classification and forced to alter their features to be recognized as valid. Formally, we can define the social burden as:
@@
Let $f$ be a classifier and $g_{f}$ a recourse algorithm. Given a sensitive group $s\in\mathcal{S}$, let $\delta((x,S=s),g_{f}(x))$ be the cost of applying the recourse provided by $g_{f}$. Then, the expected social burden $B_{f,g}^{s}$ can be defined as:
[…]
with the expectation over $P(X\mid S=s,Y=1,f(X)=0)$, and $P(f(X)=1\mid S=s,Y=1)$ indicates the true positive rate (TPR) for the given sensitive group.
Similarly to Eqn. 2, the instances obtaining a positive prediction do not count towards the overall burden. Eqn. 3 underscores the importance of considering the predictive performance of the base classifier when evaluating algorithmic recourse, a factor often overlooked in the recourse literature but central to algorithmic fairness; especially, since numerous studies have shown that machine learning classifiers tend to exhibit unequal error rates (e.g., TPR) across sensitive groups (pessach2023algorithmic). When error rates differ across groups, the burden of recourse is also likely to be unevenly distributed, potentially amplifying existing disparities. However, it is important to note that satisfying equality of opportunity (i.e., parity in TPR) does not guarantee a small gap in social burden.
@@ 3.3 Gap-Based Metrics can Mask Unfairness
Unfairness in recourse can be quantified by measuring gaps in cost or social burden across sensitive groups: $\Delta C_{f,g}=\max_{s,s^{\prime}\in\mathcal{S}}|C^{s}_{f,g}-C^{s^{\prime}}_{f,g}|$ or $\Delta B_{f,g}=\max_{s,s^{\prime}\in\mathcal{S}}|B^{s}_{f,g}-B^{s^{\prime}}_{f,g}|$. According to Equations 2 and 3, these disparities stem from three main sources. For cost gaps, the causes include: (a) disparities in AR, as captured by statistical parity (dwork2012fairness); (b) differences in the cost function, which may arise from unequal opportunities or socio-economic conditions; and (c) variations in the distribution $P(X\mid S,f(X)=0)$, reflecting potential population skews. For social burden gaps, they arise from: (a) disparities in true positive rates, captured by equality of opportunity (hardt2016equality), (b) unequal cost functions, and (c) skews in the distribution of qualified individuals denied recourse $P(X\mid S,Y=1,f(X)=0)$. Importantly, this illustrates that satisfying equality of opportunity (statistical parity) does not guarantee an equal distribution of the social burden (recourse costs) across groups.
Although comparing the gap of statistical measures across groups is a common approach in the fairness literature, it can obscure meaningful underlying unfairness. In fairness for prediction, a well-known limitation of parity-based definitions is that parity can be achieved by degrading performance for the privileged group without improving outcomes for the unprivileged one (martinez2020minimax; diana2021minimax). As a result, groups may exhibit similar expected values for a given metric, even when those values are far from optimal. In recourse, this issue is especially critical when evaluating social burden, which ideally should be as low as possible (preferably zero). A small gap in burden does not necessarily indicate fairness, as parity might be reached by increasing the burden for the privileged group rather than reducing it for the unprivileged one. However, in many real-world contexts, such as access to essential goods and services, sacrificing overall well-being is not acceptable. Therefore, fairness efforts should aim to reduce the burden for all groups, rather than simply equalizing it.
We propose drawing inspiration from approaches in the fairness-in-prediction literature that address this limitation and complement difference-based disparity measures with a Rawlsian (i.e., minimax) perspective (binns2018fairness). For instance, in addition to evaluating the gap in social burden, we can examine the worst-group social burden defined as $\max_{s\in\mathcal{S}}B_{f,g}^{s}$. This perspective’s benefits are twofold: (a) it helps identify the most vulnerable group, and (b) it serves as a base for the development of fairness-enhancing recourse frameworks that go beyond minimizing burden disparities, aiming instead to reduce the social burden for all groups.
@@
5:   $b_{f,g}^{i}\leftarrow\delta(x^{i},g_{f}(x^{i}))\mathds{1}\{y^{i}=1\}$ //Compute burden
6:   $\mathcal{Q}\leftarrow\mathcal{Q}\cup b_{f,g}^{i}$
7:  end for
8:  $f^{(t)}\leftarrow\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}\phi(i,\mathcal{Q},\alpha)\cdot\ell(f(x^{i}),y^{i})$
9:end for
10:return $f^{(T)}$
Table 2: Empirical average test results on the Adult dataset across different sensitive group characterizations. The best results for each setting are highlighted in bold, and we report whether the target metric should increase ($\uparrow$}) or decrease ($\downarrow$}). For MISOB, we set $\alpha=0.3$ in all experiments (cf. Eqn. 4). Each cell shows the average and standard deviation over 10 runs.
Sensitive Recourse Fairness acc Burden (Eq. 3) TPR Cost (Eq. 2) AR
@@
Attribute(s) Method strategy overall $\uparrow$} worst $\downarrow$} $\Delta$ $|\downarrow|$} worst $\uparrow$} $\Delta$ $|\downarrow|$} worst $\downarrow$} $\Delta$ $|\downarrow|$} worst $\uparrow$} $\Delta$ $|\downarrow|$}
Race GS - 0.81±0.02 4.56±0.01 0.03±0.02 0.27±0.03 0.08±0.08 115.69±1.19 28.37±1.95 0.06±0.01 0.06±0.02
@@
(a) WT - Race
(b) WT - Gender
(c) GS - Race
(d) GS - Gender
Figure 2: Empirical evaluation of MISOB for different $\alpha\in\{0.1,\ldots,1.0\}$. We report the average results and standard deviation (shaded area) over 10 runs, with WT (wachter2017counterfactual) and GS (laugel2017inverse), on the Adult dataset with race and gender as the sensitive attributes. The worst burden and cost (Eqn. 7) are in log scale. In brief, increasing $\alpha$, which favors high-burden instances, improves fairness guarantees, but large $\alpha$ values may hurt overall accuracy.
4 MISOB: MInimax SOcial Burden
In this section, we introduce a simple yet effective iterative training procedure, MISOB (MInimax SOcial Burden), that empirically reduces social burden (Eqn. 3) without degrading predictive performance and without requiring access to sensitive group attributes during training or inference. The method is outlined in Algorithm 1. Given a pre-trained classifier (line 1) [^5] Empirically, we observed that pretraining a base classifier to achieve reasonable accuracy before applying MISOB contributes to a more stable optimization process., the key idea is to incorporate social burden estimates directly into the training loop. At each iteration, we compute the social burden for each instance (line 5), and we save the result within $\mathcal{Q}$ (line 6). The model is then updated by minimizing a (weighted) loss function $\ell(f(x^{i}),y^{i})$ (line 8), where weights are determined by the estimated burden $\phi(i,\mathcal{Q})$. Specifically, we propose assigning weights to instances $x^{i}$ based on the following scheme:
$\phi(i,\mathcal{Q},\alpha)=1+\alpha N\dfrac{b_{f,g_{f}}^{i}}{\sum_{j\in|\mathcal{Q}|}b_{f,g_{f}}^{j}}\mathds{1}\{\beta>0\}$ (4)
where $N$ is the dataset (or batch) size, $\beta=\sum_{j\in|\mathcal{Q}|}b_{f,g_{f}}^{j}$ is the total burden with $b_{f,g}^{i}=\delta(x^{i},g_{f}(x^{i}))\mathds{1}\{y^{i}=1\}$ being the burden of instance $i$, and $\alpha$ is a hyperparameter controlling the influence of the burden term. Under this weighting scheme, each instance $x^{i}$ is assigned a weight proportional to its contribution to the total burden across the dataset (or batch). In the degenerate case, where the total burden is zero, all instances are assigned equal weights. This reweighting scheme amplifies the influence of high-burden instances during training, encouraging the classifier to prioritize decisions that reduce unnecessary or costly adjustments imposed on individuals under the current decision rule.
The proposed framework offers several advantages that enhance its practicality and broad applicability. [^6] The runtime of Algorithm 1 is determined by three factors: (i) the number of retraining steps $T$, (ii) the dataset size $N$, and (iii) the run-time complexity of the recourse method $O(K)$. In practice, datasets are typically much larger than the number of retraining iterations, and if a simple recourse method is used, we generally have $K<<T<<N$. As a result, the overall computational complexity is $O(N^{3})$. Efficiency can be further improved through batching, parallelization, or approximating recourse computations. First, the method is agnostic to both the base classifier and the chosen recourse strategy, allowing seamless integration with a wide range of existing approaches. It also operates without access to sensitive attributes during training, avoiding legal, ethical, and practical challenges associated with collecting or using such data. Further, MISOB not only improves deployability but also naturally addresses intersectional fairness. Because group definitions are not required a-priori, intersectional subgroups can be introduced post-hoc at evaluation time, enabling fairness assessments across multiple dimensions without retraining. This flexibility makes the framework well-suited for real-world applications where fairness concerns span diverse and evolving group identities.
5 Empirical Evaluation
In this section, we evaluate the effectiveness of MISOB in reducing the social burden across sensitive groups. Refer to Appendix B for further details on the experimental setting.
Experimental Setting.
We consider the real-world dataset Adult (adult_2) with sensitive groups defined by different characterizations of the sensitive attributes (race and gender), considering also intersectional groups. For the model, we consider a simple feed-forward neural network. We have chosen this base classifier because it supports multiple different recourse methods proposed in the literature. Further experiments with a linear regression model and other datasets can be found in Appendix C. In our evaluation, we consider MISOB and POSTPRO (hardt2016equality), which enhance prediction fairness in terms of equality of opportunity. We employ POSTPRO because it is compatible with any base classifier, enabling fair comparisons with MISOB under identical predictive models. We consider three widely adopted recourse methods, selected for their benchmarking popularity, available implementations, and ease of integration in an agnostic setting: Growing Spheres (GS) (laugel2017inverse), the method by wachter2017counterfactual (WT), and the Counterfactual Conditional Heterogeneous Variational Autoencoder (CCHVAE) by pawelczyk2020learning. We use the implementations and hyperparameters from the benchmark by pawelczyk1carla, and we repeat the experiments over 10 random train/test splits. We consider the worst-group values for various metrics:
$\min_{s\in\mathcal{S}}P(f(X=x)=1\mid S=s,Y=1)\;\text{(Worst TPR)}$ (5)
$\min_{s\in\mathcal{S}}P(f(X=x)=1\mid S=s)\;\text{(Worst AR)}$ (6)
$\min_{s\in\mathcal{S}}C_{f,g}^{s}\;\text{(Worst Cost)}\qquad\min_{s\in\mathcal{S}}B_{f,g}^{s}\;\text{(Worst Burden)}$ (7)
where the recourse cost $\delta$ is quantified using the $\ell_{2}$ distance, and their maximum disparity (denoted by $\Delta$) across groups:
$\max_{s,s^{\prime}\in\mathcal{S}}|C_{f,g}^{s}-C_{f,g}^{s^{\prime}}|\qquad\max_{s,s^{\prime}\in\mathcal{S}}|B_{f,g}^{s}-B_{f,g}^{s^{\prime}}|$ (8)
$\max_{s,s^{\prime}\in\mathcal{S}}|\text{TPR}^{s}-\text{TPR}^{s^{\prime}}|\qquad\max_{s,s^{\prime}\in\mathcal{S}}|\text{AR}^{s}-\text{AR}^{s^{\prime}}|$ (9)
Importantly, our method operates entirely without access to sensitive attributes. Note that computing fairness metrics-whether for prediction or recourse-requires access to sensitive attributes, which is an inherent limitation of these metrics themselves (buyl2024inherent).
Results.